# Parameterization-driven Neural Implicit Surfaces Editing

### v2

@@ 3.5 Training Losses
where $\mathbf{c}_{\text{gt}}$ and $\mathbf{c}_{\text{pred}}$ are the ground-truth and predicted colors, respectively. To ensure the shape and appearance latent codes, $\mathbf{z}_{s}$ and $\mathbf{z}_{a}$, conform to Gaussian distributions, we incorporate a regularization term, expressed as $\mathcal{L}_{\text{code}}=\|\mathbf{z}_{s}\|_{2}+\|\mathbf{z}_{a}\|_{2}$ [33]. Additionally, since both the 3D surface $\mathcal{S}$ and the parametric domain $\mathcal{D}$ are represented as signed distance fields, we employ the Eikonal loss [60, 10], defined as $\mathcal{L}_{\text{Eik}}=(\|\nabla s\|_{2}-1)^{2}$. Putting it all together, we define the total loss function as follows:
$\begin{split}\mathcal{L}=\mathcal{L}_{\text{rgb}}+&\lambda_{1}\mathcal{L}_{\text{Eik}}+\lambda_{2}\mathcal{L}_{\text{cycle}}+\lambda_{3}\mathcal{L}_{\text{smooth}}\\ &+\lambda_{4}\mathcal{L}_{\text{Lap}}+\lambda_{5}\mathcal{L}_{\text{shd}}+\lambda_{6}\mathcal{L}_{\text{code}}.\end{split}$ (13)
@@ 4 Experiments
Datasets. For our experiments, we used a customized dataset named FS-Syn, derived from the Facescape dataset [58]. From Facescale, we selected 10 human head models and configured 30 fixed viewpoints and lighting conditions for each to obtain synthesized multi-view images. We also evaluated our method on several man-made objects from the DTU dataset [17] and the OmniObject dataset [52].
Training details and metric. All the sub-networks $F_{\text{def}}$, $F_{\text{inv-def}}$, $F_{\text{sdf}}$, $F_{\text{mat}}$, and $F_{\text{shd}}$ are MLPs. See supplementary material for other detailed network hyperparameters. We trained our network using the Adam [19] optimizer. Once the parametric domain is chosen, we first trained $F_{\text{sdf}}$ for 2,000 iterations to obtain the implicit neural representation of the parametric domain. Subsequently, we fixed $F_{\text{sdf}}$ and trained other sub-networks for 2,000 epochs driven by the loss function as Equation (13). The quality of the polycube map is measured by the angle distortion and area distortion [3, 15] $\mathcal{E}_{\text{angle}},\mathcal{E}_{\text{area}}$ as follows:
$\mathcal{E}_{\text{angle}}=\frac{\cot\alpha|a|^{2}+\cot\beta|b|^{2}+\cot\gamma|c|^{2}}{4\text{area}(\Delta_{i})},$ (14)
$\mathcal{E}_{\text{area}}=\frac{1}{2}\left(\frac{\text{area}(\Delta_{i})}{\text{area}(\hat{\Delta}_{i})}+\frac{\text{area}(\hat{\Delta}_{i})}{\text{area}(\Delta_{i})}\right),$ (15)
@@
$\mathcal{S}$ $\mathcal{D}$
Figure 5: Parameterization results of objects with non-zero genus topology and complex shape.
Results. Our neural parameterization framework enables the construction of a bijective map between objects and multiple parametric domains such as spheres, cubes, and polycubes, as shown in Figure 3 and Figure 5. Additionally, our method supports co-parameterization of multiple objects and facilitates the transfer of material and shading between them. Our neural parameterization algorithm is fully compatible with existing neural rendering pipelines. Given multi-view images as input, we can reconstruct the 3D geometry and create the texture maps simultaneously. As shown in Figure 3, we can obtain high-fidelity results without any prior information.
@@
Figure 6: Parametrization results of NeP [27]
@@
Compared with NeP [27], which is the latest neural surface editing work using neural parameterization, our method eliminates the need for prior information from UV mapping. Unlike our algorithm, NeP learns a map between surfaces and the UV plane. However, it heavily relies on an existing UV mapping as initialization, and yields rather poor texture mapping results without the UV mapping as prior (see Figure 6). Moreover, the results from NeP still exhibit significant distortions even with UV prior. Our method can effectively reduce angle distortion by minimizing Laplacian loss and reduce area distortion by choosing an appropriate parametric domain. As demonstrated in Figure 7, we calculate the angle distortion $\mathcal{E}_{\text{angle}}$ of parameterization as described in [3, 15]. We can observe that the Laplacian loss can effectively reduce angle distortion.
Refer to caption w/o $\mathcal{L}_{\text{Lap}}$    w/ $\mathcal{L}_{\text{Lap}}$
Method $\mathcal{E}_{\text{angle}}$downarrow
w/o $\mathcal{L}_{\text{Lap}}$ 1.147
w/ $\mathcal{L}_{\text{Lap}}$ 1.092
Figure 7: Ablation study. Utilization of the Laplacian loss has proven effective in reducing angle distortion ($\mathcal{E}_{\text{angle}}$), especially in regions with high curvature.
GT w/o $F_{\text{shd}}$ w/ $F_{\text{shd}}$
Figure 8: Ablation study. For scenes with significantly varying illumination, it is crucial to decompose view-dependent shading from the radiance field to further enable separate editing.
Radiance decomposition. In contrast to an entangled radiance field in NeRF [29] and VolSDF [61], we decompose the radiance into view-independent material and view-dependent shading according to Equation 9. We illustrate the decomposition results in Figure 9. We would like to emphasize that it is essential to decompose the radiance according to the view-dependent correlations for scenes with varying illumination. As shown in Figure 8 (down), it fails to reconstruct the 3D geometry when we remove the $F_{\text{shd}}$, and the visual quality drops subsequently.
GT Material Shading
Figure 9: Radiance decomposition results.
@@
Figure 10: Representing the materials as textures defined in simple parametric domain enables easy editing. Also, our co-parameterization naturally supports texture transfer between different objects.
@@
Texture editing. Both NeuMesh [57] and NeP [27] can edit the object’s texture at a pixel level. We compare the editing and rendering results with NeuMesh and NeP, as shown in Figure 10. NeuMesh requires a mask of the editing region. It also re-trains the network after each edit, which is time-consuming and cannot guarantee the stability of the training results. Our method outperforms NeuMesh in terms of the visual quality. NeP, which is also neural parameterization-based method, requires a UV prior for initialization. It often yields poor parameterization results without such prior. Even with the UV prior, its parameterization results exhibit large distortion, making the edit in the 2D domain challenging. Our method is free of prior, and can yield parameterization with low distortion, thereby facilitating pixel-level editing and material transferring without network re-training.
Shading editing. We decompose the 3D radiance field into view-independent material and view-dependent shading defined in the parametric domain. Taking advantage of this decomposition and the embedding into a simple domain, we can edit both the material and shading of the 3D surfaces in an intuitive manner. As illustrated in Figure 11, we modify the shading of one model by transferring the shading from another model under different lighting conditions, while keeping the material unchanged.
@@
Figure 11: We transfer the decomposed shading from one model to another. See also the accompanying video.
@@ 5 Conclusion
We present a novel neural algorithm for parametrizing 3D surfaces represented by neural implicit functions to a user-defined parametric domain, such as a sphere, cube, or polycube. Utilizing bi-directional deformation, our method is capable of learning a bijective mapping without relying on any prior knowledge, while controlling angle distortion through the usage of a Laplacian regularizer. Furthermore, our approach seamlessly integrates with the neural rendering pipeline, enabling the reconstruction of 3D objects from multi-view images and facilitating efficient volume rendering of modified textures and shadings — eliminating the necessity for network re-training. We demonstrated the efficacy of our method on human heads and man-made objects.
@@
Figure 12: Geometry reconstruction results of VolSDF [61] and ours. Our method is capable of editing through parameterization, accompanied by an acceptable compromise in reconstruction quality.
Our method has a few limitations that require further improvement in the future. First, computational results on the FS-Syn dataset show that the reconstruction quality of our method is slightly worse than the standard SDF-based neural rendering algorithms, such as VolSDF [61] and NeuS [47]. For example, as illustrated in Fig. 12 and the right table, We report PSNR and CD compared with VolSDF [61] on Fs-Syn dataset [58].
@@
This decline in visual quality arises from the incorporation of additional modules, such as bi-directional deformation and radiance decomposition, resulting in the cessation of updates to the SDF network within our neural parameterization pipeline. Therefore, computing a parameterization without compromising the reconstruction quality is highly desired. Second, our radiance decomposition operates under the assumptions of Lambertian reflectance and grayscale shading, which might hinder its effectiveness when dealing with intricate materials or sophisticated shadings. Third, in our current implementation, we did not optimize the runtime performance. Therefore, training our model is time consuming, requiring 3 days for the 10 human heads of the FS-Syn dataset [58]. Incorporating recent advancements such as PermutoSDF [38] and Strivec [7] has the potential to significantly reduce the training and inference time.
Supplementary Material
Appendix A Implementation Details
In our framework, the bi-deformation networks $F_{\text{def}}$ and $F_{\text{inv-def}}$, the SDF network $F_{\text{sdf}}$, the texture network $F_{\text{mat}}$, and the appearance network $F_{\text{shd}}$ are all multilayer perceptrons, consisting of 8, 8, 8, 4, and 4 layers, respectively. Each of these hidden layers has 256 neurons. The latent shape codes $z_{s}$ and the appearance codes $z_{a}$ are 128-dimensional. Within the neural rendering process, we set the frequency of position encoding for positions and viewpoints as $6$ and $4$, respectively. In each iteration, we randomly sample 1024 pixels from each input image for training. As for other parameter settings related to neural field reconstruction, we refer mainly to the work of VolSDF  [61].
@@ Appendix A Implementation Details
For static settings, we remove the warp module in the NeP [27]. Then we train NeP under two settings, one is under UV ground truth as supervision and the other is not (see Figure 13). The parameter configurations for other settings can be found in  [27].
Figure 13: We compare the rendering results with NeP without prior information from UV mapping.
@@
We derived a customized dataset from Facescape [58], as shown in Figure 14. We randomly chose 10 human heads with neutral expressions from uniformly topologized models in the FaceScape dataset. Subsequently, we illuminated the human head model by moving both the light source and the camera, generating diverse lighting effects. Following this, we captured 30 images from different angles around the human head, each accompanied by its corresponding camera parameters.
Figure 14: We create the FS-Syn dataset from the FaceScape dataset. We illuminated the human head by moving both the light source and the camera, creating varying lighting effects.
Appendix B Additional Results
We also tested our method on the DTU dataset [17] as shown in Figure  15(a). Moreover, our parameterization framework can be used to seamlessly stitch segmented texture maps together, as shown in Figure 15(b). In addition to texture and shading, our framework also can edit geometric details on original models based on parameterization results and the assigned displacement map, as illustrated in Figure 15(c).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
(c) Refer to caption Refer to caption Refer to caption
Figure 15: Additional results: (a) Visualization of normal maps, neural rendering results, and parametric domain on models from the DTU dataset [17]. (b) Textures. The textures associated with the original meshes (middle) are discontinuous and exhibit significant distortion, making them less suitable for editing purposes. In contrast, the textures resulting from our parameterization (right) are globally continuous and exhibit minimal distortions. (c) Geometric detail editing. (d) Source geometry, target appearance, and texture transfer results. Durian and Apple undergo joint training as they belong to the same category.
Appendix C Geometry Editing via Modifying the Parametric Domain $\mathcal{D}$

### v3

@@ 3.5 Training Losses
where $\mathbf{c}_{\text{gt}}$ and $\mathbf{c}_{\text{pred}}$ are the ground-truth and predicted colors, respectively. To ensure the shape and appearance latent codes, $\mathbf{z}_{s}$ and $\mathbf{z}_{a}$, conform to Gaussian distributions, we incorporate a regularization term, expressed as $\mathcal{L}_{\text{code}}=\|\mathbf{z}_{s}\|_{2}+\|\mathbf{z}_{a}\|_{2}$ [33]. Additionally, with geometry represented as signed distance fields, we employ the Eikonal loss [62, 11], defined as $\mathcal{L}_{\text{Eik}}=(\|\nabla s\|_{2}-1)^{2}$. Putting it all together, we define the total loss function as follows:
$\mathcal{L}=\mathcal{L}_{\text{rgb}}+\lambda_{1}\mathcal{L}_{\text{Eik}}+\lambda_{2}\mathcal{L}_{\text{cycle}}+\lambda_{3}\mathcal{L}_{\text{smooth}}+\lambda_{4}\mathcal{L}_{\text{Lap}}+\lambda_{5}\mathcal{L}_{\text{shd}}+\lambda_{6}\mathcal{L}_{\text{code}}.$ (17)
@@ 4 Experiments
Datasets. For our experiments, we used a customized dataset named FS-Syn, derived from the Facescape dataset [59], selecting 10 human head models and configuring 30 fixed viewpoints and lighting conditions to obtain synthesized multi-view images. We also evaluated our method on several man-made objects from the OmniObject dataset [54], and the H3DS dataset [38].
GT
Normal
$\mathcal{D}$
$\mathcal{S}$
Figure 2: Parameterization results. Human heads are co-parameterized to a sphere owing to the clear geometric resemblance. Given the diverse geometry of cars, each is parameterized to a polycube domain (bottom three rows). From top to bottom: the input image, the normal map of the reconstructed geometry $\mathcal{S}$, the parametric domain $\mathcal{D}$, and the reconstructed surface $\mathcal{S}$. See supplementary material for additional results.
Training details and metrics. All the sub-networks $F_{\text{def}}$, $F_{\text{inv-def}}$, $F_{\text{sdf}}$, $F_{\text{mat}}$, and $F_{\text{shd}}$ are MLPs. See the supplementary material for other detailed network hyperparameters. Once the parametric domain is chosen, we fixed $F_{\text{sdf}}$ and trained other sub-networks for 2,000 epochs driven by the loss function as Eq. (17). We applied the standard Marching Cubes algorithm to extract triangle meshes for the zero level-set of the source surface $\mathcal{S}$ and then used the forward map to obtain the mesh representing the parametric surface. With triangle meshes, we measured the quality of the computed parameterization by the angle distortion $\mathcal{E}_{\text{angle}}$ and area distortion [4, 16] $\mathcal{E}_{\text{area}}$ as follows:
$\mathcal{E}_{\text{angle}}=\frac{\cot\alpha|a|^{2}+\cot\beta|b|^{2}+\cot\gamma|c|^{2}}{4\text{area}(\Delta_{i})},~{}~{}~{}\mathcal{E}_{\text{area}}=\frac{1}{2}\left(\frac{\text{area}(\Delta_{i})}{\text{area}(\hat{\Delta}_{i})}+\frac{\text{area}(\hat{\Delta}_{i})}{\text{area}(\Delta_{i})}\right),$
@@
1.545/1.396 1.755/1.446 1.394/1.302 1.271/1.216 1.351/1.275
Figure 3: Results of cross-polycube parameterization on the H3DS dataset [38]. We learned a common polycube domain for all human heads with the parameter $k=3$ (i.e., the polycube is formed by 3 boxes) and then computed the polycube parameterization for all the human heads simultaneously. Since human heads share similar geometries, the computed polycube parameterizations are also highly consistent. The values below each figure represent the angle distortion and the area distortion, respectively.
Results. Our framework constructs a nearly bijective map between objects and multiple parametric domains such as spheres and polycubes, as shown in Fig. 2 and Fig. 3. It supports co-parameterization of multiple objects and facilitates material and shading transfer. Fully compatible with existing neural rendering pipelines, our neural parameterization algorithm uses multi-view images to simultaneously reconstruct 3D geometry and create texture maps. As shown in Fig. 2, we achieve high-fidelity results without prior information.
@@
Figure 4: Parametrization results of NeP [27] in comparison with our method. NeP heavily relies on the initial UV prior, leading to poor texture maps when such prior information is lacking. Our method learns texture maps from a set of multi-view images without any 3D tracked mesh or UV prior as input.
@@
Compared with NeP [27], which is the latest neural surface editing work using neural parameterization, our method eliminates the need for prior information from UV mapping. Unlike our algorithm, NeP learns a map between surfaces and the UV plane. However, it heavily relies on an existing UV mapping as initialization, and yields rather poor texture mapping results without the UV mapping as prior (see Fig.4). Moreover, the results from NeP still exhibit significant distortions even with UV prior.
@@
Figure 5: Representing the materials as textures defined in a simple parametric domain enables easy editing. Also, our co-parameterization naturally supports texture transfer between different objects.
@@
Texture editing. Our parameterization results enable flexible and pixel-level editing as shown in Fig. 5. We re-paint the extracted 2D texture map using cubemaps [10] and retrieve the modified color using the established mapping when volume rendering novel images after each editing. Both NeuMesh [57] and NeP [27] can edit the object’s texture at a pixel level. We compare the editing and rendering results with NeuMesh and NeP, as shown in Fig. 5. NeuMesh requires a mask of the editing region. It also re-trains the network after each edit, which is time-consuming and cannot guarantee the stability of the training results. Our method outperforms NeuMesh in terms of the visual quality. NeP, which is also neural parameterization-based method, requires a UV prior for initialization. It often yields poor parameterization results without such prior. Even with the UV prior, its parameterization results exhibit large distortion, making the edit in the 2D domain challenging. Our method is free of prior, and can yield parameterization with low distortion, thereby facilitating pixel-level editing and material transferring without network re-training.
GT Material Shading GT Material Shading
(a) We decompose the radiance into view-dependent shadings and view-independent materials. The shading map is visualized from the front view.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
(b) We transfer the decomposed shading from one model to another. See also the accompanying video.
Figure 6: We transfer the decomposed shading from one model to another. For further illustration, please refer to the accompanying video.
Shading editing. We decompose the 3D radiance field into view-independent material and view-dependent shading defined in the parametric domain. Taking advantage of this decomposition and the embedding into a simple domain, we can edit both the material and shading of the 3D surfaces in an intuitive manner. As illustrated in Fig. 6(b), we modify the shading of one model by transferring the shading from another model under different lighting conditions, while keeping the material unchanged.
Inverse deformation. To validate the cycle loss and illustrate inverse deformation, we sample points on the polycube parametric surface $\mathcal{D}$. Subsequently, we inversely deform these points back to the surface $\mathcal{S}$. The point cloud is visualized in the right inset, and we quantify the CD (×$\times 10^{-3}$) with respect to the ground truth. The obtained low CD results confirm the efficacy of inverse deformation and cycle loss.
Ablation studies. As shown in Fig. 7 (left), the term $\mathcal{L}_{\text{smooth}}$ helps to smooth the process of the two deformations, while $\mathcal{L}_{\text{Lap}}$ further reduces the angle distortion. Our method can effectively reduce angle distortion by minimizing Laplacian loss and reduce area distortion by choosing an appropriate parametric domain. As demonstrated in Fig. 7 (right), we calculate the angle distortion $\mathcal{E}_{\text{angle}}$ of parameterization as described in [4, 16]. We can observe that the Laplacian loss can effectively reduce angle distortion.
The choice of parametric domain. The different parametric domain influences the quality of parameterization as shown in Fig. 8 (a), which is measured by area distortions. Generally, increasing the number of cubic primitives for better resembling the geometry of the target surface effectively reduces distortions but increases the complexity of the parametric domain. Finding an optimal balance between domain complexity and parameterization quality is challenging.
Refer to captionRefer to captionRefer to captionRefer to caption
GT VolSDF 4.65 Ours 4.87 Inverse 4.89
Refer to captionRefer to captionRefer to captionRefer to caption
Input w/o $\mathcal{L}_{\text{smooth}}\&\mathcal{L}_{\text{Lap}}$ $+\mathcal{L}_{\text{smooth}}$ $+\mathcal{L}_{\text{Lap}}$
$\mathcal{E}_{\text{angle}}=1.147$    $\mathcal{E}_{\text{angle}}=\mathbf{1.092}$
w/o $\mathcal{L}_{\text{Lap}}$    w/ $\mathcal{L}_{\text{Lap}}$
Figure 7: Ablation studies. On the left, we calculate the CD (×$\times 10^{-3}$) based on the point cloud extracted from the inverse deformation module and present the ablation studies for $\mathcal{L}_{\text{smooth}}$ and $\mathcal{L}_{\text{Lap}}$. On the right, utilization of the Laplacian loss has proven effective in reducing angle distortion ($\mathcal{E}_{\text{angle}}$), especially in regions with high curvature.
@@ 5 Conclusion
We introduce a novel neural algorithm for parameterizing 3D surfaces represented by neural implicit functions to a user-defined parametric domain, such as spheres and polycubes. This parametric domain is learned given some user-specified hyperparameters. Utilizing bi-directional deformation, our method is capable of learning a nearly bijective mapping without relying on any prior knowledge, while controlling angle distortion through the usage of a Laplacian regularizer. Furthermore, our approach seamlessly integrates with the neural rendering pipeline, enabling the reconstruction of 3D objects from multi-view images and facilitating efficient volume rendering of modified textures and shadings — eliminating the necessity for network re-training. We demonstrated the efficacy of our method on human heads and man-made objects.
Our method has a few limitations that require further improvement in the future. First, quantitative comparisons on the FS-Syn dataset show that the reconstruction quality of our method is slightly worse than the standard SDF-based neural rendering algorithms, such as VolSDF [61] and NeuS [49]. For example, as illustrated in Fig. 8 (b) and, We report PSNR and CD compared with VolSDF [61] on Fs-Syn dataset [59]. This decline in visual quality arises from the incorporation of additional modules, such as bi-directional deformation and radiance decomposition, resulting in the cessation of updates to the SDF network within our neural parameterization pipeline. Therefore, computing a parameterization without compromising the reconstruction quality is highly desired. Second, our radiance decomposition operates under the assumptions of Lambertian reflectance and grayscale shading, which might hinder its effectiveness when dealing with intricate materials or sophisticated shadings.
Refer to captionRefer to captionRefer to captionRefer to caption
$\mathcal{S}$ $\mathcal{E}_{\text{area}}=2.505$ $6.094$ $1.633$
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
(a) (b)
Figure 8: We present area distortion $\mathcal{E}_{\text{area}}$ based on the different choices of the parametric domain in (a) and geometry reconstruction results of VolSDF [61] and our method in (b). Our method is capable of editing through parameterization, accompanied by an acceptable compromise in reconstruction quality.
Acknowledgment
This study is supported under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s). This project is also partially supported by the Ministry of Education, Singapore, under its Academic Research Fund Grants (MOE-T2EP20220-0005 & RT19/22).
Appendix 0.A Implementation Details and Ablation Studies
In our framework, the bi-deformation networks $F_{\text{def}}$ and $F_{\text{inv-def}}$, the material network $F_{\text{mat}}$, and the shading network $F_{\text{shd}}$ are all multilayer perceptrons, consisting of 8, 8, 4, and 4 layers, respectively. Each of these hidden layers has 256 neurons. The latent shape codes $z_{s}$ and the appearance codes $z_{a}$ are 128-dimensional. Within the neural rendering process, we set the frequency of position encoding for positions and viewpoints as $6$ and $4$, respectively. We trained our network using the Adam optimizer [18]. In each iteration, we randomly sample 1024 pixels from each input image for training. As for other parameter settings related to neural field reconstruction, we refer mainly to the work of VolSDF [61].
@@ Appendix 0.A Implementation Details and Ablation Studies
We remove the warp module in the NeP [27] for static settings. Then we train NeP under two settings, one is under UV reference as supervision and the other is not (see Fig. 9). The parameter configurations for other settings can be found in the original paper [27].
Figure 9: (Left) We compare the rendering results with NeP without prior information from UV mapping. (Right) We show texture editing results demonstrating that our method offers a more intuitive editing approach, capable of modifying the material without changing the shading.
Ablation studies on shading module.
As shown in Fig. 10 (left), a larger value ($\lambda_{\text{shd}}=0.1$) leads to broken and inconsistent shading. Conversely, omitting $\mathcal{L}_{\text{shd}}$ results in the shading module learning material intensity that is not caused by the lighting.
Refer to captionRefer to captionRefer to captionRefer to caption
$\lambda_{\text{shd}}=0.1$ $0.01$ $10.0$
Figure 10: Ablation studies on the shading module demonstrate how the learning of lighting varies with different values of $\lambda_{\text{shd}}$ (left). We created the FS-Syn dataset from the FaceScape dataset [59] by illuminating human heads while moving both the light source and the camera, thereby producing varying lighting effects (right).
@@
We derived a customized dataset from Facescape [59], as shown in Fig. 10. We randomly chose 10 human heads with neutral expressions from uniformly topologized models in the FaceScape [59] dataset. Subsequently, we illuminated the human head model by moving both the light source and the camera, generating diverse lighting effects. Following this, we captured 30 images from different angles around the human head, each accompanied by its corresponding camera parameters.
Appendix 0.B Geometry Editing via Modifying the Parametric Domain $\mathcal{D}$